OpenStudy (anonymous):

bx - cx = -c solve for x and find any restrictions what is a restriction???????? HELPPPP

jimthompson5910 (jim_thompson5910):

the idea is to isolate x bx - cx = -c x(b - c) = -c ... factor out the GCF x on the left side x = -c/(b-c) .. divide both sides by b-c to isolate x

jimthompson5910 (jim_thompson5910):

now the question is: can b and c be any real numbers? or are there restrictions?

jimthompson5910 (jim_thompson5910):

btw a restriction is a condition that limits what a variable can be for example, if you had 1/x, you can have x be any real number BUT x cannot be 0 (since you cannot divide by zero)

OpenStudy (anonymous):

OHHHHHHHHHH thanks so much this makes more sense

OpenStudy (anonymous):

@jim_thompson5910 there arnt any restrictions right??

jimthompson5910 (jim_thompson5910):

there are actually

jimthompson5910 (jim_thompson5910):

b-c CANNOT be zero if it were, then b - c = 0 b - c + c = 0+c b = c so if b = c, then b-c is 0

jimthompson5910 (jim_thompson5910):

so the restriction (the only restriction) is that b cannot be equal to c \[\large b \neq c\]

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

oh ok thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome